the inquiry club at jefferson university has compiled a book that exposes the private lives of many of the professors on campus. Economics majors in the club estimate that total revenue from sales of the books is given by the equation TR =120Q -0.1Q3

a. over what output range is demand elastic


1
Expert's answer
2021-02-12T10:47:15-0500

a. Demand becomes elastic after the point, at which TR = max or MR = 0.

"MR = TR'(Q) = 120 - 0.3 Q^2 = 0,"

"Q^2 = 400,"

Q = 20 units (Q = -20 is not possible in our case).